# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ II Overview of the 2D statistical models with long-range interaction
For the 2D LR-XY model, it is crucial to distinguish between the Berezinskii–Kosterlitz–Thouless (BKT) transition, where $\xi$ diverges exponentially as $\xi\sim\exp(b/\sqrt{t})$, and the second-order transition, where $\xi\sim t^{-\nu}$, with $t=(T-T_{c})/T_{c}$ and $\nu$ as the correlation length exponent. In our studies [cpl_42_7_070002, yao2024nonclassical], we found that for $\sigma\leq 2$, both direct observations and extrapolations of $\xi$ clearly follow a power-law divergence, confirming a genuine second-order transition, whereas for $\sigma>2$, an exponential scaling of $\xi$ is demonstrated, illustrating a BKT transition. A similar pattern emerges for the Heisenberg case [2dlrheisen]. For $\sigma\leq 2$, $\xi$ exhibits a power-law divergence near some finite $T_{c}$ as, $\xi\sim t^{-\nu}$. For $\sigma>2$, the scaling $\xi\sim\exp(b/T)$ is identified, where $b$ is a non-universal constant, which indicates the absence of finite-temperature transitions, consistent with asymptotic freedom in the 2D SR Heisenberg model.
Finally, from the viewpoint of critical exponents, our results are also consistent with the above observations. Both $\eta$ and $\nu$ exhibit a smooth evolution across the nonclassical regime. For small $\sigma$, $\eta$ follows the long-range prediction $\eta=2-\sigma$. However, as $\sigma$ approaches 2, $\eta$ decreases more slowly than $2-\sigma$ and shows a slight discontinuity at $\sigma=2$. For $\sigma>2$, $\eta$ approaches and saturates to the short-range value $\eta_{\mathrm{SR}}$. This deviation near $\sigma=2$ provides another quantitative signature of the crossover in universality. Meanwhile, field-theoretical analyses [giachetti2021, giachetti2022] predict an additional QLRO phase in the XY case for $7/4<\sigma\leq 2$, but such an intermediate phase was not observed in our numerical studies [cpl_42_7_070002, yao2024nonclassical].
@@
To address this challenge, we consider the $q$-state Potts model [potts1952some] whose Hamiltonian is given by
@@
where $s_{i}=1,2,\dots,q$ is the Potts spin, and $\delta_{s_{i},s_{j}}$ is the Kronecker delta, equal to $1$ if $s_{i}=s_{j}$ and 0 otherwise. For the 2D LR case, we set $K_{ij}=c(\sigma,L)/r_{ij}^{2+\sigma}$, similar to Eq. (1). Notably, this model can be mapped to the FK representation [fortuin1972random, swendsen1987], where the configuration consists of clusters, and its partition function can be written as
@@
For the 2D $q$-state Potts model [Jacobsen_2014], one can define a critical polynomial as
@@
where $\langle\mathcal{R}_{2}\rangle$ denotes the probability that a spanning cluster emerges to connect itself by wrapping around the periodic boundaries along both the $x$ and $y$ directions, and $\langle\mathcal{R}_{0}\rangle$ represents the probability that no cluster wraps at all. At the critical point, both $\langle\mathcal{R}_{2}\rangle$ and $\langle\mathcal{R}_{0}\rangle$ take universal values, but $R_{p}=0$ provides a more convenient criterion. Owing to the duality of planar graphs [Jacobsen_2014], this exact relation $R_{p}=0$ holds at the critical point of the SR model, without any finite-size corrections, for the $q$-state Potts model on the square or triangular lattices. Although an exact proof is absent for non-planar graphs, previous numerical studies [xu2021, PhysRevE.98.062101] have shown that $R_{p}=0$ remains valid for models belonging to the SR universality class. Therefore, deviations of $R_{p}$ from zero can serve as a sensitive probe for detecting the LR–SR crossover.
For the percolation case ($q=1$), in addition to the conventional method of generating percolation configurations for a given $K$, we also employ the event-based ensemble method [li_explosive_2023]. This method enables simulations directly at pseudo-critical points, thereby avoiding potential systematic errors arising from inaccurate critical-point determination. Utilizing both the conventional and event-based methods, we achieved system sizes up to $L=16384$ and determined the universal values of various dimensionless quantities ($Q_{m}$ and $R_{p}$) and the critical exponents, including $\eta$, $y_{t}=1/\nu$, and the shortest-path exponent $d_{\rm min}$ [2dlrperco], with high precision. The results reveal a pronounced jump in both the universal ratios and the critical exponents at $\sigma=2$, providing compelling evidence that the crossover from SR to LR universality occurs exactly at $\sigma=2$, in contradiction to Sak’s criterion.
@@
Finally, it is worth noting the 2D LR-UF model. Similar to percolation, this is a purely graphical model but with a continuous symmetry [PhysRevLett.93.080601]. Despite its simplicity, the model exhibits rich critical behavior: for $\sigma>2$, no finite-temperature phase transition occurs, while for $\sigma\leq 2$, a second-order phase transition appears. The resulting phase diagram closely resembles that of the 2D LR-Heisenberg model.
@@ III Results for 2D Long-range Ising Model
2.5 0.369 446(2) 0.001(2) 0.857(1)      0.860(3)[Luijten2002] 0.250(1)      0.246(8)[Luijten2002] 1.007(5)
NN 0.440 686…[PhysRev.65.117] [Jacobsen_2014] 0.856 216(1)[Kamieniarz_1993] 0.25 [PhysRev.65.117] [PhysRev.65.117]
In this section, we present our simulation results for the 2D LR-Ising model, which provide evidence that the universality boundary between the LR and SR regimes lies at $\sigma_{*}=2$. To achieve this, we employ an enhanced variant of the Luijten–Blöte cluster algorithm [luijten1995, luijten1997, michel2019], enabling simulations up to $L=8192$. We extract the critical points and universal values from the critical polynomial $R_{p}$ and the Binder ratio $Q_{m}$, and obtain the critical exponents from the susceptibility $\chi$ and the scaled covariance $g^{(x)}_{ER}$ (defined later in Eq. (13)). By comparing with previous works [Luijten2002, angelini2014, horita2017], we find that our numerical results are more precise with smaller errors, clearly demonstrating the location of $\sigma_{*}=2$. Our results are summarized in Table 1.
@@ III.1 Critical points
It is worth noting that Horita et al. [horita2017] also adopted a normalized formulation for the LR-Ising model, allowing a direct comparison of the estimated critical points. After converting their estimations to our normalization conventions, the critical points are fully consistent: Horita et al. [horita2017] reported $\beta_{c}=0.329\,134(8)$ for $\sigma=1.75$ and $\beta_{c}=0.344\,445(11)$ for $\sigma=2$, while our results are $\beta_{c}=0.329\,136(1)$ and $\beta_{c}=0.344\,439(2)$, respectively. These values agree within the error bars, with our simulations offering significantly higher precision.
@@ III.2 Dimensionless quantities
As shown in Fig. A1 in Appendix A, one can already see with significant confidence that in the limit of large system size $L\to\infty$, the universal values for $\sigma=1.75$ and $1.875$ deviate from those of the SR case. Specifically, the universal value of $R_{p}$ shifts significantly away from zero [Jacobsen_2014], while $Q_{m}$ clearly departs from $0.856216(1)$ [blote1995]. At $\sigma=2$, a noticeable deviation is still observed for $R_{p}$, whereas for $Q_{m}$, the difference remains within the error bars and is therefore hardly distinguishable.
@@
To provide an overall perspective, we summarize in Fig. 5 the extrapolated critical values of $R_{p}$ and $Q_{m}$ as functions of $\sigma$. For $Q_{m}$, we also include a comparison with previous studies [Luijten2002, horita2017]. As $\sigma$ decreases from the SR to the LR regime, $R_{p}$ shows a clear discontinuity at $\sigma=2$, followed by a gradual decline for $\sigma<2$. This discontinuity is particularly pronounced in $R_{p}$, whereas for $Q_{m}$ it is less evident due to larger statistical uncertainties. Compared with previous results, our extrapolated values of $Q_{m}$ fall well within their reported error bars. However, while our results clearly reveal a universality change as $\sigma$ decreases from $2$, the previous studies could hardly distinguish such a transition because of the large uncertainties. The extrapolated universal values for $R_{p}$ and $Q_{m}$ are summarized in Table 1.
Figure 5: Extrapolated universal values of $R_{p}$ and $Q_{m}$ as functions of $\sigma$. The black lines denote the SR universal values ($R_{p}=0$, $Q_{m}=0.856216$). A clear jump is observed at $\sigma=2$ for $R_{p}$, marking the crossover from SR to LR universality. For $\sigma<2$, both quantities continuously decrease from the SR values, while for $\sigma>2$, they remain consistent with the SR values. Previous studies [Luijten2002, horita2017] are also shown for comparison.
Finally, we remark on the approach of Ref. [horita2017], which introduced a nonlinear “self-combined” Binder ratio,
@@ III.3 Critical exponents
Critical exponents provide a direct characterization of universality, and consequently, much of the existing literature on this issue has focused on their estimation [Luijten2002, picco2012, angelini2014]. However, obtaining reliable exponents is challenging because the precise location of the critical point is often uncertain, and strong finite-size effects can obscure the true asymptotic behavior. Here, based on critical points determined with high precision above and simulations of large system sizes up to $L=8192$, we achieve a more accurate extraction of the exponents. Moreover, the availability of numerous previous studies enables a direct and meaningful comparison of our results with earlier findings.
@@
Moreover, from the perspective of the geometric FK representation, the fraction of the largest-cluster size $C_{1}$ over the whole lattice can also act as an order parameter, similar to the magnetization density. Namely, the size of the largest cluster scales as $C_{1}\sim L^{2-\eta/2}$ in 2D. We plot $C_{1}/L^{2}$ versus $L$ on a log-log scale, and similarly, it is clearly seen that when $\sigma\geq 2$, the slope $\eta/2$ deviates from the SR case. In short, with sufficiently large system sizes in a wide range from $L=512$ to $L=8192$, the bare-eye view of the finite-size scaling of $\chi$ and $C_{1}$ can already yield strong evidence that $\sigma_{*}=2$. In contrast, the data for $L\leq 512$ suffer from strong finite-size corrections (see insets of Fig. 6), and, to extract correct values of $\eta$, these corrections should be carefully and systematically taken into account. Note that the maximum system sizes are $L=1000$ in Ref [Luijten2002] and $5120$ in Ref [picco2012].
@@
In Fig. 7(a), we plot our estimates of $\eta$ as a function of $\sigma$. According to Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm SR})$, where $\eta_{\rm SR}=1/4$ for the 2D Ising model, as shown by the dashed line. Our estimates of $\eta$ are plotted as red dots. It is clear that there is a deviation from Sak’s criterion for $\sigma\in[1.75,2]$. Notably, at $\sigma=2$, $\eta$ is already larger than $1/4$, which is consistent with the universal value of $R_{p}$ discussed in the previous section. Moreover, other previous numerical results are also plotted here. The green dots from Luijten et al. [Luijten2002], which have large errors in this regime, cannot clearly distinguish the deviation; specifically, their central values are actually close to our results, but the errors are very large. As for the blue dots from Picco [picco2012], they are very close to our results, and his estimates, extending to $\sigma=1.6$, also imply that the deviation already starts near $\sigma=1.6$.
@@
where $\mathcal{E}=-J_{\rm nn}\sum_{\langle i,j\rangle}S_{i}S_{j}$ denotes the NN energy density, and $J_{\rm nn}=c(\sigma,L)$ is the NN coupling strength. Here, $\mathcal{R}^{(x)}=1$ if at least one cluster wraps in the $x$ direction (otherwise $\mathcal{R}^{(x)}=0$). This covariance scales as $g^{(x)}_{ER}\sim L^{y_{t}}$ [PhysRevE.99.042150], which can be used to extract the exponent $y_{t}$.
@@
Figure 7: The value of critical exponents ($\eta$ and $y_{t}$) at various $\sigma$. The solid lines denote the SR values for $\eta$ and $y_{t}$. (a) The red dots denote the estimates of $\eta$ in this work. Luijten and Picco’s results are also contained [Luijten2002, picco2012], respectively denoted by green and blue dots. The dotted lines correspond to Sak’s prediction: $\eta=\text{max}(2-\eta,\frac{1}{4})$. Obviously, our results deviate from Sak’s prediction, suggesting that for $1.75<\sigma\leq 2$, there exists a discontinuous jump at $\sigma=2$, indicating that the short-range behavior is not recovered. (b) The red dots denote the estimates of $y_{t}$ in this work. The dotted line represents the results from second-order $\epsilon$ expansion [fisher1972] (note that $y_{t}=1/\nu$).
In Fig. 7(b), we plot the estimates of $y_{t}$ as functions of $\sigma$. The blue and purple dots at $\sigma=1.6$ correspond to the results from Picco [picco2012] and Angelini et al. [angelini2014], respectively. The dashed line represents the second-order $\epsilon$ expansion from Fisher et al. [fisher1972]. We can see that $y_{t}$ remains close to $1$ across the investigated range, making it less effective as an indicator of universality. The deviations from the $\epsilon$ expansion are also quite clear, which again imply the limitation of the mean-field expansion, similar to the case of $\eta$.

### v2

@@ II Overview of the 2D statistical models with long-range interaction
For the 2D LR-XY model, it is crucial to distinguish between the Berezinskii–Kosterlitz–Thouless (BKT) transition, where $\xi$ diverges exponentially as $\xi\sim\exp(b/\sqrt{t})$, and the second-order transition, where $\xi\sim t^{-\nu}$, with $t=(T-T_{c})/T_{c}$ and $\nu$ as the correlation length exponent. In our studies [54, 60], we found that for $\sigma\leq 2$, both direct observations and extrapolations of $\xi$ clearly follow a power-law divergence, confirming a genuine second-order transition, whereas for $\sigma>2$, an exponential scaling of $\xi$ is demonstrated, illustrating a BKT transition. A similar pattern emerges for the Heisenberg case [59]. For $\sigma\leq 2$, $\xi$ exhibits a power-law divergence near some finite $T_{c}$ as, $\xi\sim t^{-\nu}$. For $\sigma>2$, the scaling $\xi\sim\exp(b/T)$ is identified, where $b$ is a non-universal constant, which indicates the absence of finite-temperature transitions, consistent with asymptotic freedom in the 2D SR Heisenberg model.
Finally, from the viewpoint of critical exponents, our results are also consistent with the above observations. Both $\eta$ and $\nu$ exhibit a smooth evolution across the nonclassical regime. For small $\sigma$, $\eta$ follows the long-range prediction $\eta=2-\sigma$. However, as $\sigma$ approaches 2, $\eta$ decreases more slowly than $2-\sigma$ and shows a slight discontinuity at $\sigma=2$. For $\sigma>2$, $\eta$ approaches and saturates to the short-range value $\eta_{\mathrm{SR}}$. This deviation near $\sigma=2$ provides another quantitative signature of the crossover in universality. Meanwhile, field-theoretical analyses [22, 23] predict an additional QLRO phase in the XY case for $7/4<\sigma\leq 2$, but such an intermediate phase was not observed in our numerical studies [54, 60].
@@
To address this challenge, we consider the $q$-state Potts model [47] whose Hamiltonian is given by
@@
where $s_{i}=1,2,\dots,q$ is the Potts spin, and $\delta_{s_{i},s_{j}}$ is the Kronecker delta, equal to $1$ if $s_{i}=s_{j}$ and 0 otherwise. For the 2D LR case, we set $K_{ij}=c(\sigma,L)/r_{ij}^{2+\sigma}$, similar to Eq. (1). Notably, this model can be mapped to the FK representation [19, 51], where the configuration consists of clusters, and its partition function can be written as
@@
For the 2D $q$-state Potts model [27], one can define a critical polynomial as
@@
where $\langle\mathcal{R}_{2}\rangle$ denotes the probability that a spanning cluster emerges to connect itself by wrapping around the periodic boundaries along both the $x$ and $y$ directions, and $\langle\mathcal{R}_{0}\rangle$ represents the probability that no cluster wraps at all. At the critical point, both $\langle\mathcal{R}_{2}\rangle$ and $\langle\mathcal{R}_{0}\rangle$ take universal values, but $R_{p}=0$ provides a more convenient criterion. Owing to the duality of planar graphs [27], this exact relation $R_{p}=0$ holds at the critical point of the NN model, without any finite-size corrections, for the $q$-state Potts model on the square or triangular lattices. Although an exact proof is absent for non-planar graphs, previous numerical studies [55, 44] have shown that $R_{p}=0$ remains valid for models belonging to the SR universality class. Therefore, deviations of $R_{p}$ from zero can serve as a sensitive probe for detecting the LR–SR crossover.
For the percolation case ($q=1$), in addition to the conventional method of generating percolation configurations for a given $K$, we also employ the event-based ensemble method [35]. This method enables simulations directly at pseudo-critical points, thereby avoiding potential systematic errors arising from inaccurate critical-point determination. Utilizing both the conventional and event-based methods, we achieved system sizes up to $L=16384$ and determined the universal values of various dimensionless quantities ($Q_{m}$ and $R_{p}$) and the critical exponents, including $\eta$, $y_{t}=1/\nu$, and the shortest-path exponent $d_{\rm min}$ [36], with high precision. The results reveal a pronounced jump in both the universal ratios and the critical exponents at $\sigma=2$, providing compelling evidence that the crossover from SR to LR universality occurs exactly at $\sigma=2$, in contradiction to Sak’s criterion.
@@
Finally, it is worth noting the 2D LR-UF model. Similar to percolation, this is a purely graphical model but with a continuous symmetry [13]. Despite its simplicity, the model exhibits rich critical behavior: for $\sigma>2$, no finite-temperature phase transition occurs, while for $\sigma\leq 2$, a second-order phase transition appears. The resulting phase diagram closely resembles that of the 2D LR-Heisenberg model.
@@ III Results for 2D Long-range Ising Model
2.5 0.369 446(2) 0.001(2) 0.857(1)      0.860(3)[39] 0.250(1)      0.246(8)[39] 1.007(5)
NN 0.440 686…[43] [27] 0.856 216(1)[30] 0.25 [43] [43]
In this section, we present our simulation results for the 2D LR-Ising model, which provide evidence that the universality boundary between the LR and SR regimes lies at $\sigma_{*}=2$. To achieve this, we employ an enhanced variant of the Luijten–Blöte cluster algorithm [40, 38, 42], enabling simulations up to $L=8192$. We extract the critical points and universal values from the critical polynomial $R_{p}$ and the Binder ratio $Q_{m}$, and obtain the critical exponents from the susceptibility $\chi$ and the scaled covariance $g^{(x)}_{ER}$ (defined later in Eq. (13)). By comparing with previous works [39, 5, 25], we find that our numerical results are more precise with smaller errors, clearly demonstrating the location of $\sigma_{*}=2$. Our results are summarized in Table 1.
@@ III.1 Critical points
It is worth noting that Horita et al. [25] also adopted a normalized formulation for the LR-Ising model, allowing a direct comparison of the estimated critical points. After converting their estimations to our normalization conventions, the critical points are fully consistent: Horita et al. [25] reported $\beta_{c}=0.329\,134(8)$ for $\sigma=1.75$ and $\beta_{c}=0.344\,445(11)$ for $\sigma=2$, while our results are $\beta_{c}=0.329\,136(1)$ and $\beta_{c}=0.344\,439(2)$, respectively. These values agree within the error bars, with our simulations offering significantly higher precision.
@@ III.2 Dimensionless quantities
As shown in Fig. A1 in Appendix A, one can already see with significant confidence that in the limit of large system size $L\to\infty$, the universal values for $\sigma=1.75$ and $1.875$ deviate from those of the SR case. Specifically, the universal value of $R_{p}$ shifts significantly away from zero [27], while $Q_{m}$ clearly departs from $0.856216(1)$ [7]. At $\sigma=2$, a noticeable deviation is still observed for $R_{p}$, whereas for $Q_{m}$, the difference remains within the error bars and is therefore hardly distinguishable.
@@
To provide an overall perspective, we summarize in Fig. 5 the extrapolated critical values of $R_{p}$ and $Q_{m}$ as functions of $\sigma$. For $Q_{m}$, we also include a comparison with previous studies [39, 25]. As $\sigma$ decreases from the SR to the LR regime, $R_{p}$ shows a clear discontinuity at $\sigma=2$, followed by a gradual decline for $\sigma<2$. This discontinuity is particularly pronounced in $R_{p}$, whereas for $Q_{m}$ it is less evident due to larger statistical uncertainties. Compared with previous results, our extrapolated values of $Q_{m}$ fall well within their reported error bars. However, while our results clearly reveal a universality change as $\sigma$ decreases from $2$, the previous studies could hardly distinguish such a transition because of the large uncertainties. The extrapolated universal values for $R_{p}$ and $Q_{m}$ are summarized in Table 1.
Figure 5: Extrapolated universal values of $R_{p}$ and $Q_{m}$ as functions of $\sigma$. The black lines denote the SR universal values ($R_{p}=0$, $Q_{m}=0.856216$). A clear jump is observed at $\sigma=2$ for $R_{p}$, marking the crossover from SR to LR universality. For $\sigma<2$, both quantities continuously decrease from the SR values, while for $\sigma>2$, they remain consistent with the SR values. Previous studies [39, 25] are also shown for comparison.
Finally, we remark on the approach of Ref. [25], which introduced a nonlinear “self-combined” Binder ratio,
@@ III.3 Critical exponents
Critical exponents provide a direct characterization of universality, and consequently, much of the existing literature on this issue has focused on their estimation [39, 46, 5]. However, obtaining reliable exponents is challenging because the precise location of the critical point is often uncertain, and strong finite-size effects can obscure the true asymptotic behavior. Here, based on critical points determined with high precision above and simulations of large system sizes up to $L=8192$, we achieve a more accurate extraction of the exponents. Moreover, the availability of numerous previous studies enables a direct and meaningful comparison of our results with earlier findings.
@@
Moreover, from the perspective of the geometric FK representation, the fraction of the largest-cluster size $C_{1}$ over the whole lattice can also act as an order parameter, similar to the magnetization density. Namely, the size of the largest cluster scales as $C_{1}\sim L^{2-\eta/2}$ in 2D. We plot $C_{1}/L^{2}$ versus $L$ on a log-log scale, and similarly, it is clearly seen that when $\sigma\geq 2$, the slope $\eta/2$ deviates from the SR case. In short, with sufficiently large system sizes in a wide range from $L=512$ to $L=8192$, the bare-eye view of the finite-size scaling of $\chi$ and $C_{1}$ can already yield strong evidence that $\sigma_{*}=2$. In contrast, the data for $L\leq 512$ suffer from strong finite-size corrections (see insets of Fig. 6), and, to extract correct values of $\eta$, these corrections should be carefully and systematically taken into account. Note that the maximum system sizes are $L=1000$ in Ref [39] and $5120$ in Ref [46].
@@
In Fig. 7(a), we plot our estimates of $\eta$ as a function of $\sigma$. According to Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm SR})$, where $\eta_{\rm SR}=1/4$ for the 2D Ising model, as shown by the dashed line. Our estimates of $\eta$ are plotted as red dots. It is clear that there is a deviation from Sak’s criterion for $\sigma\in[1.75,2]$. Notably, at $\sigma=2$, $\eta$ is already larger than $1/4$, which is consistent with the universal value of $R_{p}$ discussed in the previous section. Moreover, other previous numerical results are also plotted here. The green dots from Luijten et al. [39], which have large errors in this regime, cannot clearly distinguish the deviation; specifically, their central values are actually close to our results, but the errors are very large. As for the blue dots from Picco [46], they are very close to our results, and his estimates, extending to $\sigma=1.6$, also imply that the deviation already starts near $\sigma=1.6$.
@@
where $\mathcal{E}=-J_{\rm nn}\sum_{\langle i,j\rangle}S_{i}S_{j}$ denotes the NN energy density, and $J_{\rm nn}=c(\sigma,L)$ is the NN coupling strength. Here, $\mathcal{R}^{(x)}=1$ if at least one cluster wraps in the $x$ direction (otherwise $\mathcal{R}^{(x)}=0$). This covariance scales as $g^{(x)}_{ER}\sim L^{y_{t}}$ [26], which can be used to extract the exponent $y_{t}$.
@@
Figure 7: The value of critical exponents ($\eta$ and $y_{t}$) at various $\sigma$. The solid lines denote the SR values for $\eta$ and $y_{t}$. (a) The red dots denote the estimates of $\eta$ in this work. Luijten and Picco’s results are also contained [39, 46], respectively denoted by green and blue dots. The dotted lines correspond to Sak’s prediction: $\eta=\text{max}(2-\eta,\frac{1}{4})$. Obviously, our results deviate from Sak’s prediction, suggesting that for $1.75<\sigma\leq 2$, there exists a discontinuous jump at $\sigma=2$, indicating that the short-range behavior is not recovered. (b) The red dots denote the estimates of $y_{t}$ in this work. The dotted line represents the results from second-order $\epsilon$ expansion [18] (note that $y_{t}=1/\nu$).
In Fig. 7(b), we plot the estimates of $y_{t}$ as functions of $\sigma$. The blue and purple dots at $\sigma=1.6$ correspond to the results from Picco [46] and Angelini et al. [5], respectively. The dashed line represents the second-order $\epsilon$ expansion from Fisher et al. [18]. We can see that $y_{t}$ remains close to $1$ across the investigated range, making it less effective as an indicator of universality. The deviations from the $\epsilon$ expansion are also quite clear, which again imply the limitation of the mean-field expansion, similar to the case of $\eta$.